Considering the encircle part of the pumping system as shown below. Assuming a fresh water flowing in the discharge pipe with a velocity of 5 m/s.


1
Expert's answer
2022-02-28T16:08:13-0500

"v_1S_1=v_2S_2,"

"v_2=v_1\frac{S_1}{S_2}=5\cdot\frac{12}6=10~\frac ms."